ON THE PROFINITE TOPOLOGY ON DOUBLES OF GROUPS

RITA GITIK

Abstract. We give a sufficient condition for a finitely generated subgroup of a double of a group to be closed in the profinite topology. It follows that quasiconvex subgroups in doubles of certain negatively curved LERF groups are closed in the profinite topology. In particular, certain doubles are LERF if they have the finitely generated intersection property.